Mathematics

writhe the following fraction in decimal form.

 ( a) 24/100

(b) 24/10

(C)24/1000
Answer #1

(a) 24/100=0.24

(b)24/10 = 2.4

(c)24/1000 =0.024


answered by: anonymous